# Performance Analysis of NR Sidelink and Wi-Fi Coexistence Networks in Unlicensed Spectrum

### v1

@@ VI Conclusion
In this paper, we investigate methods to enhance the overall performance of coexistence systems operating in the unlicensed spectrum while ensuring fair coexistence between SL-U users and Wi-Fi users. First, we propose a CCHA mechanism that integrates channel access with resource allocation. This mechanism enhances spectrum fairness in coexistence systems while improving spectrum utilization. Next, we formulate a joint mathematical model for channel access and power control, aiming to achieve a balance between fairness and transmission rate. Finally, we develop a C-GHDRL algorithm framework. The framework enables joint optimization of channel access and power control through collaborative interactions between the BS and SL-U users, thereby significantly enhancing the overall performance of the coexistence system. Future research will focus on exploring more efficient collaborative mechanisms under conditions of incomplete global information.

### v2

@@ VI Conclusion
In this paper, we investigate methods to enhance the overall performance of coexistence systems operating in the unlicensed spectrum while ensuring fair coexistence between SL-U users and Wi-Fi users. First, we propose a CCHA mechanism that integrates channel access with resource allocation. This mechanism enhances spectrum fairness in coexistence systems while improving spectrum utilization. Next, we develop a Cooperative Sub-goal-based Hierarchical Deep Reinforcement Learning (C-GHDRL) framework to further optimize spectrum utilization. The framework enables joint optimization of channel access and power control through cooperative interaction between the BS and SL-U users, thereby significantly improving the overall performance of the coexistence system. Finally, we formulate a joint mathematical model for channel access and power control, aiming to balance fairness and transmission rate by defining an appropriate reward function within C-GHDRL algorithm. Future research will focus on exploring more efficient collaborative mechanisms under conditions of incomplete global information.